Acme Company is considering replacing outdated production equipment that will allow for production cost savings of $20,000 per month. The new equipment will have a five-year life and cost $800,000, with an estimated salvage value of $50,000. Acme's cost of capital is 10%.

Calculate the payback period and the accounting rate of return for the new production equipment.
Answer #1

Payback period = Initial investment/Annual Cost saving = 800000/240000 = 3.33 years

Accounting rate of return = Net income/Average investment = 90000/425000 = 21.18%